$SPAD/src/input oct.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool oct.output
)set message test on
)set message auto off
)clear all
--S 1 of 15
oci1 := octon(1,2,3,4,5,6,7,8)
--R
--R
--R   (1)  1 + 2i + 3j + 4k + 5E + 6I + 7J + 8K
--R                                                      Type: Octonion(Integer)
--E 1
--S 2 of 15
oci2 := octon(7,2,3,-4,5,6,-7,0)
--R
--R
--R   (2)  7 + 2i + 3j - 4k + 5E + 6I - 7J
--R                                                      Type: Octonion(Integer)
--E 2
--S 3 of 15
oci3 := octon(quatern(-7,-12,3,-10), quatern(5,6,9,0))
--R
--R
--R   (3)  - 7 - 12i + 3j - 10k + 5E + 6I + 9J
--R                                                      Type: Octonion(Integer)
--E 3
--S 4 of 15
(oci1 * oci2) * oci3 - oci1 * (oci2 * oci3)
--R
--R
--R   (4)  2696i - 2928j - 4072k + 16E - 1192I + 832J + 2616K
--R                                                      Type: Octonion(Integer)
--E 4
--S 5 of 15
[real oci1, imagi oci1, imagj oci1, imagk oci1, imagE oci1, _
 imagI oci1, imagJ oci1, imagK oci1]

--R
--R
--R   (5)  [1,2,3,4,5,6,7,8]
--R                                                  Type: List(PositiveInteger)
--E 5
--S 6 of 15
q : Quaternion Polynomial Integer := quatern(q1, qi, qj, qk)
--R
--R
--R   (6)  q1 + qi i + qj j + qk k
--R                                        Type: Quaternion(Polynomial(Integer))
--E 6
--S 7 of 15
E : Octonion Polynomial Integer:= octon(0,0,0,0,1,0,0,0)
--R
--R
--R   (7)  E
--R                                          Type: Octonion(Polynomial(Integer))
--E 7
--S 8 of 15
q * E
--R
--R
--R   (8)  q1 E + qi I + qj J + qk K
--R                                          Type: Octonion(Polynomial(Integer))
--E 8
--S 9 of 15
E * q
--R
--R
--R   (9)  q1 E - qi I - qj J - qk K
--R                                          Type: Octonion(Polynomial(Integer))
--E 9
--S 10 of 15
q * 1$(Octonion Polynomial Integer)
--R
--R
--R   (10)  q1 + qi i + qj j + qk k
--R                                          Type: Octonion(Polynomial(Integer))
--E 10
--S 11 of 15
1$(Octonion Polynomial Integer) * q
--R
--R

--R   (11)  q1 + qi i + qj j + qk k
--R                                          Type: Octonion(Polynomial(Integer))
--E 11
--S 12 of 15
o : Octonion Polynomial Integer := octon(o1, oi, oj, ok, oE, oI, oJ, oK)
--R
--R
--R   (12)  o1 + oi i + oj j + ok k + oE E + oI I + oJ J + oK K
--R                                          Type: Octonion(Polynomial(Integer))
--E 12
--S 13 of 15
norm o
--R
--R
--R           2     2     2     2     2     2     2     2
--R   (13)  ok  + oj  + oi  + oK  + oJ  + oI  + oE  + o1
--R                                                    Type: Polynomial(Integer)
--E 13
--S 14 of 15
p : Octonion Polynomial Integer := octon(p1, pi, pj, pk, pE, pI, pJ, pK)
--R
--R
--R   (14)  p1 + pi i + pj j + pk k + pE E + pI I + pJ J + pK K
--R                                          Type: Octonion(Polynomial(Integer))
--E 14
--S 15 of 15
norm(o*p)-norm(p)*norm(p)
--R
--R
--R   (15)
--R         4
--R     - pk
--R   +
--R              2      2      2      2      2      2      2     2     2     2
--R         - 2pj  - 2pi  - 2pK  - 2pJ  - 2pI  - 2pE  - 2p1  + ok  + oj  + oi
--R       +
--R           2     2     2     2     2
--R         oK  + oJ  + oI  + oE  + o1
--R    *
--R         2
--R       pk
--R   +
--R         4
--R     - pj
--R   +
--R              2      2      2      2      2      2     2     2     2     2     2

--R         - 2pi  - 2pK  - 2pJ  - 2pI  - 2pE  - 2p1  + ok  + oj  + oi  + oK  + oJ
--R       +
--R           2     2     2
--R         oI  + oE  + o1
--R    *
--R         2
--R       pj
--R   +
--R         4
--R     - pi
--R   +
--R              2      2      2      2      2     2     2     2     2     2     2
--R         - 2pK  - 2pJ  - 2pI  - 2pE  - 2p1  + ok  + oj  + oi  + oK  + oJ  + oI
--R       +
--R           2     2
--R         oE  + o1
--R    *
--R         2
--R       pi
--R   +
--R         4
--R     - pK
--R   +
--R              2      2      2      2     2     2     2     2     2     2     2
--R         - 2pJ  - 2pI  - 2pE  - 2p1  + ok  + oj  + oi  + oK  + oJ  + oI  + oE
--R       +
--R           2
--R         o1
--R    *
--R         2
--R       pK
--R   +
--R         4
--R     - pJ
--R   +
--R           2      2      2     2     2     2     2     2     2     2     2   2
--R     (- 2pI  - 2pE  - 2p1  + ok  + oj  + oi  + oK  + oJ  + oI  + oE  + o1 )pJ
--R   +
--R         4         2      2     2     2     2     2     2     2     2     2   2
--R     - pI  + (- 2pE  - 2p1  + ok  + oj  + oi  + oK  + oJ  + oI  + oE  + o1 )pI
--R   +
--R         4         2     2     2     2     2     2     2     2     2   2     4
--R     - pE  + (- 2p1  + ok  + oj  + oi  + oK  + oJ  + oI  + oE  + o1 )pE  - p1
--R   +
--R        2     2     2     2     2     2     2     2   2
--R     (ok  + oj  + oi  + oK  + oJ  + oI  + oE  + o1 )p1
--R                                                    Type: Polynomial(Integer)
--E 15
)spool
)lisp (bye)

———- 7

References^{[1] nothing}

8